Carson​ Electronics' management has long viewed BGT Electronics as an industry leader and uses this firm as a model firm for analyzing its own performance. The balance sheets and income statements for the two firms are found​ here:

A. Calculate the following ratios for both Carson and​ BGT:

Current ratio                                      Times interest earned

Inventory turnover                         Total asset turnover

Operating profit margin                                Operating return on assets

Debt ratio                                            Average collection period

Fixed asset turnover                      Return on equity

B. Analyze the differences you observe between the two firms. Comment on what you view as weaknesses in the performance of Carson as compared to BGT that​ Carson's management might focus on to improve its operations.

Carson​ Electronics, Inc.

Balance Sheet​ ($000)

BGT​ Electronics, Inc.

Balance Sheet​ ($000)

Cash

$1,950

$1,510

Accounts receivable

4,510

6,030

Inventories

1,480

2,510

Current assets

$7,940

$10,050

Net fixed assets

15,970

25,030

Total assets

$23,910

$35,080

Accounts payable

$2,550

$4,970

Accrued expenses

960

1,510

​Short-term notes payable

3,480

1,540

Current liabilities

$6,990

$8,020

​Long-term debt

8,020

4,050

​Owners' equity

8,900

23,010

Total liabilities and​ owners' equity

$23,910

$35,080

Carson​ Electronics, Inc.

Income Statement​ ($000)

BGT​ Electronics, Inc.

Income Statement​ ($000)

Net sales​ (all credit)

$47,980

$70,000

Cost of goods sold

(36,050)

(42,010)

Gross profit

$11,930

$27,990

Operating expenses

(7,960)

(11,990)

Net operating income

$3,970

$16,000

Interest expense

(1,130)

(600)

Earnings before taxes

$2,840

$15,400

Income taxes

​(35 %35%​)

( 994 )

(5,390)

Net income

$ 1, 846

$10,010
Answer #1

(A) :

(i) Current Ratio = Current assets/Current liabilities = (Cash + Accounts receivable + Inventory)/(Accounts payable + Accrued expenses + Short term notes payable)

Carson: Current Ratio = ($1,950 + $4,510 + $1,480)/($2,550 + $960 + $3,480) = $7,940/$6,990 = 1.13

BGT : Current Ratio = ($1,510 + $6,030 + $2,510)/($4,970 + $1,510 + $1,540) = $10,050/$8,020 = 1.25

(ii) Times interest earned = Earnings before interest and tax/Interest

Carson : Times interest earned = $3,970/$1,130 = 3.51

BGT: Times interest earned = $16,000/$600 = 26.67

(iii) Inventory turnover = Cost of goods sold/Inventory

Carson: Inventory turnover = $36,050/$1,480 = 24.36

BGT : Inventory turnover = $42,010/$2,510 = 16.74

(iv) Total Assets turnover = Sales/Total assets

Carson : Total assets turnover = $47,980/23,910 = 2.01

BGT: Total assets turnover = $70,000/$35,080 = 1.99

(v) Operating profit margin = (Operating profit/Sales)*100

Carson: Operating profit margin = ($3,970/$47,980)*100 = 8.27%

BGT: Operating profit margin = ($16,000/$70,000)*100 = 22.86%

(vi) Operating return on assets = (Operating return/Total assets) *100

Carson: Operating return on assets = ($3,970/$23,910)*100 = 16.6%

BGT: Operating return on assets = ($16,000/$35,080)*100 = 45.61%

(vii) Debt Ratio = Total liabilities/Total assets = (Current liabilities + Long term debt) /Total assets

Carson: Debt ratio = ($6,990 + $8,020)/$23,910 = 0.63

BGT = ($8,020 + $4,050)/$35,080 = 0.34

(viii) Average collection period = (Accounts receivable/Sales)*365 days

Carson: Average collection period = ($4,510/$47,980)*365 = 34.3 days

BGT: Average collection period = ($6,030/$70,000)*365 = 31.4 days

(ix) Fixed assets turnover = Sales/Fixed Assets

Carson: Fixed assets turnover = $47,980/$15,970 = 3

BGT : Fixed assets turnover = $70,000/$25,030 = 2.8

(x) Return on equity = Profit after tax/Net worth = Profit after tax/(Equity + Reserves)

Carson : Return on equity = $1,846/($8,900 + $0) = 0.21

BGT: Return on equity = $10,010/($23,010 + $0) = 0.43

(B) (i) Times interest earned ratio of Carson Electronics is much lower than BGT Electronics and it affects the profit margin.

(ii) Operating Return on assets and operating profit margin of Carson Electronics is much lower than BGT Electronics and it shows Carson Electronics's inefficiency.

(iii) Debt ratio of Carson Electronics is much higher than BGT Electronics and it results in lower profit and higher interest payment.

(iv) Return on equity of Carson Electronics is lower than BGT Electronics and it shows Carson Electronics's weaker performance.